# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v2

@@ Abstract.
This paper investigates neuron dropout as a post-processing bias mitigation for deep neural networks (DNNs). Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications. While neural networks are exceptionally good at finding statistical patterns from data, they may encode and amplify existing biases from the historical data. Existing bias mitigation algorithms often require modifying the input dataset or the learning algorithms. We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve the fairness of pre-trained DNNs. However, finding the ideal set of neurons to drop is a combinatorial problem.
We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference after training. Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility. We show that our design of randomized algorithms is effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation. We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of search algorithms on the results. Finally, we empirically and conceptually compare NeuFair to different state-of-the-art bias mitigators.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austria
@@ 1. Introduction
Broadly, fairness mitigation can be applied in pre-processing stage (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing step (e.g., changing the loss function to include fairness constraints during training), or post-processing stage (e.g., changing the logic of a pre-trained model). However, when the decision logic of AI is encoded via deep neural networks, it becomes challenging to mitigate unfairness due to its black-box uninterpretable nature.
@@
However, finding the optimal subset of dropout neurons is a combinatorial problem that requires an exhaustive search in the subset of all combinations of neuron dropouts, and hence is intractable. To overcome the computational challenges, we explore the class of randomized algorithms with Markov chain Monte Carlo (MCMC) strategies to efficiently explore the search space with statistical guarantees. In doing so, we pose the following research questions:
RQ1. How successful are randomized algorithms in fairness repairing of DNNs via dropouts?
@@
We evaluate NeuFair over $7$ deep neural network benchmarks trained over $5$ socially critical applications with significant fairness implications. We found that NeuFair can improve fairness (up to 69%) with minimal utility degradation in most cases. We also report a pathological case and reasons behind a failure of NeuFair. We also observe that NeuFair can simultaneously improve fairness and utility and provide intuitive explanations of such phenomena. Furthermore, we examine different hyperparameter configuration options of randomized algorithms. While some hyperparameters always influence fairness with positive or negative impacts; we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark. Finally, we show the effectiveness of the SA algorithm, compared to the RW and the state-of-the-art post-processing (bias) mitigator (Monjezi et al., 2023).
@@
    Experimental evaluations. We implement the randomized algorithms in NeuFair and evaluate its effectiveness and efficiency vis-a-vis the state-of-the-art techniques.
@@ 2. Background
We provide a background on the model utility and fairness measures. Then, we briefly discuss various bias mitigation methods from the AI fairness literature.
@@ 2.1. Notions of Model Utility
Accuracy can be used to gauge the overall performance of a classifier. However, for imbalanced datasets used in fairness evaluations, accuracy is a poor metric as the number of negative samples far outweighs the positive samples. For example, predicting all samples as negative (0 true positives) in the Bank (Dua and Graff, 2017a) dataset yields an accuracy of 88% but the F1 score would be 0/undefined. Since F1 is defined as the harmonic mean of precision and recall, rather than the arithmetic mean, it penalizes performance significantly if either precision or recall is low.
@@ 2.2. Notions of Model Fairness
Equal Opportunity is a relaxed variant of Equalized Odds with $Y=1$ (Hardt et al., 2016). In other words, Equalized Odds require the true positive and false positive rates to be equal across all sensitive groups. In contrast, Equal Opportunity only requires the true positive rate to be equal across all sensitive groups.
@@ 3.1. Syntax and Semantics of DNN
A deep neural network (DNN) encodes a function $\mathcal{D}:X\to[0,1]^{2}$ where $X$ consists of the set of protected attributes $X_{1}\times X_{2}\cdots\times X_{m}$ and non-protected attributes $X_{m+1}\cdots\times Y_{m+r}$. The DNN model is parameterized by the input dimension $m{+}r$, the output dimension $2$, the depth of hidden layers $n+1$, and the weights of its hidden layers $W_{0},W_{1},\ldots,W_{n}$. We describe the hidden layers with $\mathcal{M}\leftarrow[L_{I},L_{0},\ldots,L_{n},L_{O}]$, where $L_{I}$ and $L_{O}$ are the input and output layers, respectively, and ${L_{i}},\forall i\in[0,n]$, is the hidden layers. We assume that there exists a subset of neurons $\mathcal{N}\in{L_{i}},\forall i\in[0,n]$ in the hidden layers that disparately contribute to unfairness.
@@ 3.2. Inference Time Dropout for Fairness
Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of DNNs by preventing overfitting. Dropout sets all $w_{i}$ to 0 for a random set of neurons in the hidden layers with some probability during training. Once training is complete, dropout is not used, and all the neurons in the DNN are utilized to make predictions. While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the DNN during inference after training can significantly improve fairness with a minimal impact on model performance. However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
We consider a binary vector as the neuron state with $s=\{0,1\}^{N}$, where $N=\sum_{i=0}^{n}|L_{i}|$, $s_{i}$ indicates whether the neuron $i$ is dropped or not and $n$ is the number of layers. A pre-trained DNN model $\mathcal{D}$ does not include any dropouts, and hence all the indicators are 00.
@@ 1. RQ1
    How successful are randomized algorithms in fairness repairing of DNNs via dropouts?
@@ 5.1. Datasets and Models
We evaluate NeuFair with five different datasets from fairness literature. For two of the datasets, we consider two different protected groups, which effectively results in a total of 7 different benchmarks. Table 5.3 presents of an overview of the datasets and protected groups. The Adult Census Income (Dua and Graff, 2017b), Bank Marketing (Dua and Graff, 2017a), Compas Software (ProPublica, 2021), Default Credit  (Dua and Graff, 2017c), and Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) are binary classification tasks to predict whether an individual has income over $50$K, is likely to subscribe, has a low reoffending risk, is likely to default on the credit card payment, and is likely to utilize medical benefits, respectively.
We used DNNs trained over the dataset benchmarks as the machine learning model. Table 1 highlights the architectures of the DNN models used for each dataset. We use ReLU as the activation function after the linear layers. For the Bank, Default, and Compas datasets we use a dropout of 0.2 during training. For the Adult and MEPS16 datasets, we set the dropout to 0.1. For the Compas dataset, we use Adam as the optimizer with a learning rate of $0.001$. For the other datasets, we use SGD with a learning rate of $0.01$. For data preprocessing, we utilize standard techniques such as one-hot encoding for categorical features followed by min-max or standard scaling for numerical features. For the Compas dataset, we use a version of the dataset that has 12 features after feature selection that was used in (Tizpaz-Niari et al., 2022). The modified Compas dataset is available online [^4] https://github.com/Tizpaz/Parfait-ML/blob/main/subjects/datasets/compas.
@@ 5.2. Technical Details
We run all our experiments on a desktop running Ubuntu 22.04.3 LTS with an Intel(R) Core(TM) i7-7700 CPU @ 3.60GHz processor, 32GB RAM, and a 1 TB HDD. The neural networks and repair algorithms were implemented using python3.8, torch==2.0.1,
@@ 5.3. Experimental Setup
$C(s)=EOD_{s}+3.0\cdot EOD_{s_{0}}\cdot\mathbbm{1}(F1_{s}<(0.98\cdot F1_{s_{0}})).$
The F1 threshold is 98% of the validation F1 score of the unfair DNN, i.e., we tolerate a 2% degradation in the F1 score to improve fairness. A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold. For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons depending on the overall size of the network. The $n_{u}$ values are 50, 24, 24, 20, and 135 for Adult, Compas, Bank, Default, and MEPS16 respectively. For all datasets, we use a train/validation/test split of 60%/20%/20%. We set the threshold for the acceptance probabilities in the temperature initialization algorithm (Ben-Ameur, 2004) to $0.75$.
@@ 5.3.1. Effectiveness of randomized algorithms on mitigating unfairness using dropout (RQ1)
Figure 2 and Figure 3 highlight the cost of the best state found for the SA and RW over time, respectively, for 10 different seeds per data subject using the logarithmic scale. In all cases, we see a positive trend where the cost improves over time, and the search identifies better subsets of neurons to drop. We note the seemingly flat plots correspond to those cases where a good state was found in the beginning. The exception is the black line in the sub-plot for the Default dataset in both figures as it corresponds to a failed run; i.e., the best state found has an F1 score lesser than 0.98×0.98$\times$0.98 × the F1 score of the initial state. Out of 70 runs (10 per benchmark), the SA and RW found a desirable state with 0.98×0.98$\times$0.98 × the initial F1 score in 69 cases. In addition, we re-run the experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$. The progress graphs of the RW strategy are “steep” initially and as the initial states are unfair and are heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold. However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability. So, the progress graphs are not as steep initially.
@@
Pathological Case. Table 5 highlights the EOD scores for COMPAS (Sex) with the original and repaired networks. Here, we observe that SA improves the EOD in 4 cases, does not improve in 1 case, and has a worse EOD in 5 cases. RW improves the EOD in 5 cases and has a worse EOD in the other 5 cases. There are two reasons for the different trends in COMPAS (Sex). First, there is a discrepancy in the original model statistics between the validation and test splits. The difference arises because the DNN tends to overfit slightly on the training/validation sets. This is evident from model metrics in Table 3. Second, the COMPAS datasets by default are quite fair (according to EOD) compared to other datasets as their unfairness is around 2%. During the search, the SA algorithm correctly identifies a desirable state according to the validation dataset and performs better than RW. However, given the low initial unfairness and the model’s tendency to perform better on the validation/training sets, we observe minor discrepancies in the test set. The average unfairness, however, increases by 0.399% (2.522% to 2.921%), which is negligible. For both NeuFair algorithms, the validation fairness improves in the COMPAS (Sex) experiment.
@@ 5.3.2. Dropout strategies improving both fairness and utility (RQ2)
The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is the negative class and 1 is positive) as neurons continue to be dropped out from the DNN. The increase in negative predictions favors the true negatives as the datasets are highly imbalanced with more negative samples. The Default, Bank, and MEPS16 datasets have 78%, 88%, and 83% of the data belonging to the negative class. The trained DNN models were optimized for F1 score on the validation dataset to improve performance on the underrepresented positive class. The accuracy increases as the increase in true negatives outweighs the drop in true positives. The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and false negatives increase.
@@ 5.3.4. Comparing to the state-of-the-art (RQ4)
We compare the efficacy of NeuFair against a state-of-the-art post-processing bias mitigator. Dice (Monjezi et al., 2023) used a fault localization technique via do logic to pinpoint a single neuron that significantly influences fairness. The approach computes the amounts of variation in the outcomes of each layer that depends on the sensitive attributes and de-activates a single neuron with the highest effects on fairness. Since the search space evaluates a single neuron, we brute-force the entire search space of one neuron dropout in a linear time. We note that the best improvement, feasible by Dice, cannot be more than the brute-force search. Table 6 highlights the test fairness after dropping a single neuron that most impacts fairness. The results show that NeuFair substantially outperforms DICE in improving fairness. We also note a similar trend with COMPAS (Sex), where the unfairness increases slightly. The largest discrepancy in fairness improvement between NeuFair and Dice is observed in the MEPS16 dataset, with an absolute difference of 10.778% and a relative improvement of 56.124% (19.204% and 8.426%) using the SA algorithm.

### v3

@@ Abstract.
This paper investigates neuron dropout as a post-processing bias mitigation method for deep neural networks (DNNs). Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications. While DNNs are exceptional at learning statistical patterns from data, they may encode and amplify historical biases. Existing bias mitigation algorithms often require modifying the input dataset or the learning algorithms. We posit that prevalent dropout methods may be an effective and less intrusive approach to improve fairness of pre-trained DNNs during inference. However, finding the ideal set of neurons to drop is a combinatorial problem.
We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference. Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility. We show that NeuFair is efficient and effective in improving fairness (up to 69%) with minimal or no model performance degradation. We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of NeuFair on the results. Finally, we empirically and conceptually compare NeuFair to different state-of-the-art bias mitigators.
Machine Learning, Bias Mitigation, AI Ethics
copyright: acmlicenseddoi: 10.1145/3650212.3680380journalyear: 2024isbn: 979-8-4007-0612-7/24/09conference: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis; September 16–20, 2024; Vienna, Austriabooktitle: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA ’24), September 16–20, 2024, Vienna, Austriasubmissionid: issta24main-p1278-pccs: Software and its engineering Search-based software engineeringccs: Computing methodologies Machine learning
@@ 1. Introduction
Broadly, fairness mitigation can be applied in the pre-processing (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing (e.g., changing the loss function to include fairness constraints during training), or post-processing (e.g., changing the logic of a pre-trained model) stage. However, when the decision logic of AI is encoded via DNNs, it becomes challenging to mitigate unfairness due to its black-box uninterpretable nature.
@@
However, finding the optimal subset of dropout neurons is an intractable combinatorial problem that requires an exhaustive search over all possible combinations of neuron dropouts. To overcome the computational challenges, we explore the class of randomized algorithms with Markov chain Monte Carlo (MCMC) strategies to efficiently explore the search space with statistical guarantees. In doing so, we pose the following research questions:
RQ1. How successful are randomized algorithms in repairing the unfairness of DNNs via dropouts?
@@
We evaluate NeuFair over $7$ deep neural network benchmarks trained over $5$ socially critical applications with significant fairness implications. We found that NeuFair can improve fairness (up to 69%) with minimal utility degradation in most cases. We also report a pathological case and reasons behind a failure of NeuFair. We also observe that NeuFair can simultaneously improve fairness and utility and provide intuitive explanations of such phenomena. Furthermore, we examine different hyperparameter configuration options of randomized algorithms. While some hyperparameters always influence fairness with positive or negative impacts, we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark. Finally, we show the effectiveness of the SA algorithm, compared to the RW and the state-of-the-art post-processing (bias) mitigator (Monjezi et al., 2023).
@@
    Experimental evaluations. We implement the randomized algorithms in NeuFair and evaluate their effectiveness and efficiency vis-a-vis the state-of-the-art techniques.
@@ 2. Background
In this section, we provide a background on the various model utility and fairness metrics.
@@ 2.1. Notions of Model Utility
Accuracy can be used to gauge the overall performance of a classifier. However, accuracy is a poor metric for imbalanced datasets commonly used in fairness evaluations as the number of negative samples far outweighs the positive samples. For example, predicting all samples as negative (0 true positives) in the Bank (Dua and Graff, 2017a) dataset yields an accuracy of 88% but the F1 score would be 0 or undefined. Since F1 is defined as the harmonic mean of precision and recall, rather than the arithmetic mean, it penalizes performance significantly if either precision or recall is low.
@@ 2.2. Notions of Model Fairness
Equal Opportunity is a relaxed variant of Equalized Odds with $Y=1$ (Hardt et al., 2016). Equalized Odds require the true positive and false positive rates to be equal across all sensitive groups. However, Equal Opportunity only requires the true positive rate to be equal across all sensitive groups.
@@ 3.1. Syntax and Semantics of DNN
A deep neural network (DNN) encodes a function $\mathcal{D}:X\to[0,1]^{2}$ where $X$ consists of the set of protected attributes $X_{1}\times X_{2}\cdots\times X_{m}$ and non-protected attributes $X_{m+1}\cdots\times Y_{m+r}$. The DNN model is parameterized by the input dimension $m{+}r$, the output dimension $2$, the depth of hidden layers $n+1$, and the weights of its hidden layers $W_{0},W_{1},\ldots,W_{n}$. We describe the hidden layers with $\mathcal{M}\leftarrow[L_{I},L_{0},\ldots,L_{n},L_{O}]$, where $L_{I}$ and $L_{O}$ are the input and output layers, respectively, and ${L_{i}},\forall i\in[0,n]$, are the hidden layers. We assume that there exists a subset of neurons $\mathcal{N}\in{L_{i}},\forall i\in[0,n]$ in the hidden layers that disparately contribute to unfairness.
@@ 3.2. Inference Time Dropout for Fairness
Dropout (Srivastava et al., 2014) is a technique proposed to improve the performance of DNNs by preventing overfitting. Dropout sets all $w_{i}$ to 0 for a random set of neurons in the hidden layers with some probability during training. Once training is complete, dropout is not used, and all the neurons in the DNN are utilized to make predictions. While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the DNN during inference after training can significantly improve fairness with a minimal impact on performance. However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
We consider a binary vector as the neuron state with $s=\{0,1\}^{N}$, where $N=\sum_{i=0}^{n}|L_{i}|$ and $s_{i}$ indicates whether the neuron $i$ is dropped or not and $n$ is the number of layers. A pre-trained DNN model $\mathcal{D}$ does not include any dropouts; hence, all the indicators are 00.
@@ 1. RQ1
    How successful are randomized algorithms in repairing the unfairness of DNNs via dropouts?
@@ 5.1. Datasets and Models
We evaluate NeuFair with five different datasets from fairness literature. For two of the datasets, we consider two different protected groups, which effectively results in a total of 7 benchmarks. Table 5.3 presents an overview of the datasets and protected groups. The Adult Census Income (Dua and Graff, 2017b), Bank Marketing (Dua and Graff, 2017a), Compas Software (ProPublica, 2021), Default Credit  (Dua and Graff, 2017c), and Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) are binary classification tasks to predict whether an individual has income over $50$K, is likely to subscribe, has a low reoffending risk, is likely to default on the credit card payment, and is likely to utilize medical benefits, respectively.
We used DNNs trained over the dataset benchmarks as the machine learning model. Table 1 highlights the architectures of the DNN models used for each dataset. We use ReLU as the activation function after the linear layers. For the Bank, Default, and Compas datasets we use a dropout of 0.2 during training. For the Adult and MEPS16 datasets, we set the dropout to 0.1. For the Compas dataset, we use Adam as the optimizer with a learning rate of $0.001$. For the other datasets, we use SGD with a learning rate of $0.01$. For data preprocessing, we utilize standard techniques such as one-hot encoding for categorical features followed by min-max or standard scaling for numerical features. For the Compas dataset, we use a version of the dataset used in (Tizpaz-Niari et al., 2022) that has 12 features after feature selection. The modified Compas dataset is available online [^4] https://github.com/Tizpaz/Parfait-ML/blob/main/subjects/datasets/compas.
@@ 5.2. Technical Details
[…]
@@ 5.3. Experimental Setup
$C(s)=EOD_{s}+3.0\cdot EOD_{s_{0}}\cdot\mathds{1}(F1_{s}<(0.98\cdot F1_{s_{0}})).$
The F1 threshold is 98% of the validation F1 score of the unfair DNN, i.e., we tolerate a 2% degradation in the F1 score to improve fairness. A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold. For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons. The $n_{u}$ values are 50, 24, 24, 20, and 135 for Adult, Compas, Bank, Default, and MEPS16 respectively. For all datasets, we use a train/validation/test split of 60%/20%/20%. We set the threshold for the acceptance probabilities in the temperature initialization algorithm (Ben-Ameur, 2004) to $0.75$.
@@ 5.3.1. Effectiveness of randomized algorithms on mitigating unfairness using dropout (RQ1)
Figure 2 and Figure 3 highlight the cost of the best state found for the SA and RW over time, respectively, for 10 different seeds per data subject using the logarithmic scale. In all cases, we see a positive trend where the cost improves over time, and the search identifies better subsets of neurons to drop. We note the seemingly flat plots correspond to those cases where a good state was found in the beginning. The exception is the black line in the sub-plot for the Default dataset in both figures as it corresponds to a failed run; i.e., the best state found has an F1 score lesser than 0.98×0.98$\times$0.98 × the F1 score of the initial state. Out of 70 runs (10 per benchmark), the SA and RW found a desirable state with 0.98×0.98$\times$0.98 × the initial F1 score in 69 cases. In addition, we re-run the failed Default experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$. The progress graphs of the RW strategy are “steep” initially as the initial states are unfair and are heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold. However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability. Thus, the progress graphs are not as steep initially.
@@
Pathological Case. Table 5 highlights the individual EOD scores for all seeds of COMPAS (Sex) with the original and repaired networks. Here, we observe that SA improves the EOD in 4 cases, does not improve in 1 case, and has a worse EOD in 5 cases. RW improves the EOD in 5 cases and has a worse EOD in the other 5 cases. There are two reasons for the different trends in COMPAS (Sex). First, there is a discrepancy in the original model statistics between the validation and test splits. The difference arises because the DNN tends to overfit slightly on the training and validation sets. This is evident from model metrics in Table 3. Second, the COMPAS datasets by default are quite fair (according to EOD) compared to other datasets as their unfairness is around 2%. During the search, the SA algorithm correctly identifies a desirable state according to the validation dataset and performs better than RW. However, given the low initial unfairness and the model’s tendency to perform better on the training and validation sets, we observe minor discrepancies in the test set. The average unfairness, however, increases by 0.399% (2.522% to 2.921%), which is negligible. For both NeuFair algorithms, the validation fairness improves in the COMPAS (Sex) experiment.
@@ 5.3.2. Dropout strategies improving both fairness and utility (RQ2)
The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is the negative class and 1 is positive) as neurons continue to be dropped out from the DNN. The increase in negative predictions favors the true negatives as the datasets are highly imbalanced with more negative samples. The Default, Bank, and MEPS16 datasets have 78%, 88%, and 83% of the data belonging to the negative class. The trained DNN models were optimized for F1 score on the validation dataset to improve performance on the underrepresented positive class. The accuracy increases as the increase in true negatives far outweighs the drop in true positives. The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and the false negatives increase.
@@ 5.3.4. Comparing to the state-of-the-art (RQ4)
We compare the efficacy of NeuFair against a state-of-the-art post-processing bias mitigator. Dice (Monjezi et al., 2023) uses a fault localization technique via do logic to pinpoint a single neuron that significantly influences fairness. The approach computes the amounts of variation in the outcomes of each layer that depends on the sensitive attributes and de-activates a single neuron with the highest effects on fairness. Since the search space evaluates a single neuron, we brute-force the entire search space of one neuron dropout in linear time. We note that the best improvement, feasible by Dice, cannot be more than the brute-force search. Table 6 highlights the test fairness after dropping a single neuron that most impacts fairness. The results show that NeuFair substantially outperforms DICE in improving fairness. We also note a similar trend with COMPAS (Sex), where the unfairness increases slightly. The largest discrepancy in fairness improvement between NeuFair and Dice is observed in the MEPS16 dataset, with an absolute difference of 10.778% and a relative improvement of 56.124% (19.204% and 8.426%) using SA.